### v1

# Sequential model confidence sets

Abstract
In most prediction and estimation situations, scientists consider various statistical models for the same problem, and naturally want to select amongst the best. Hansen et al. (2011) provide a powerful solution to this problem by the so-called model confidence set, a subset of the original set of available models that contains the best models with a given level of confidence. Importantly, model confidence sets respect the underlying selection uncertainty by being flexible in size. However, they presuppose a fixed sample size which stands in contrast to the fact that model selection and forecast evaluation are inherently sequential tasks where we successively collect new data and where the decision to continue or conclude a study may depend on the previous outcomes. In this article, we extend model confidence sets sequentially over time by relying on sequential testing methods. Recently, e-processes and confidence sequences have been introduced as new, safe methods for assessing statistical evidence. Sequential model confidence sets allow to continuously monitor the models’ performances and come with time-uniform, nonasymptotic coverage guarantees.
Keywords: Model confidence set, forecast evaluation, model selection, sequential inference, multiple testing.
1 Introduction
In science as well as in our daily life, we frequently encounter situations in which multiple statistical models or forecasts are available for the same problem and where we have to decide which model(s) or forecast(s) we want to trust. We might think for example of multiple inflation regression models with respect to different sets of covariates or of multiple weather institutions which issue precipitation predictions for the next day. In such situations, we naturally want to select the best model(s) or forecast(s), where the term “best” is defined in terms of a user-specified criterion.
The model confidence set (MCS) proposed by Hansen et al. (2011) provides a promising solution to this problem. It departs from standard practice where just a single model is selected according to some appropriate loss and nothing is said about the uncertainty associated with this selection. The MCS takes this uncertainty into account by reducing the original set of available models to a smaller set of flexible size that contains the best models with a given level of confidence. Model confidence sets are of great importance in applications without an obvious benchmark, may be easily constructed by the MCS algorithm, and are widely applied in the econometrics community; see, e.g., Weron (2014), Corbet et al. (2019) or Masini et al. (2023) amongst others. However, even though the MCS provides an appealing solution to a highly important problem, it comes with limitations. In particular, it is assumed that the sequences of observed loss differences are stationary, an assumption which may be questionable in practise since we may expect that the models evolve over time, and correct for errors and systematic biases by using past information. Apart from relying on strong stationarity assumptions, the MCS requires a sample of some fixed size chosen independently of the data. In other words, we assume that we observe losses of the models over some prespecified evaluation period in order to compute the MCS only once at the end of the period. This procedure highly contrasts our natural urge to assess the models on a regular basis by successively including past observations to the available dataset. Consider for example different weather institutions that predict the accumulated precipitation for the following day on a daily-basis. Then, firstly, it is not given that the initially better models remain superior until the end of our study. Secondly, we want to assess the forecasters and monitor their performance sequentially, say, at the end of each day, week or month, since the forecasters issue their forecasts sequentially as well and may behave nonstationary. If we gather evidence that one institution is worse than the others but the evidence is not yet statistically significant, then we want to continue to collect evidence for this hypothesis without loosing the information of the earlier observations. Crucially, we want to decide at the end of each day, week or month whether to continue or conclude the study, depending on the previous outcomes.
In summary, model selection and forecast evaluation are inherently sequential tasks and we should seek for methods that allow to sequentially choose amongst the best models while respecting the underlying selection uncertainty. In this article, we address this need and contribute with sequential model confidence sets. Sequential model confidence sets generalize the core idea of Hansen et al. (2011) sequentially over time and enjoy time-uniform coverage guarantees of the theoretically superior objects. Our methods rely on e-processes and time-uniform confidence sequences, which lay the foundation for safe anytime-valid inference, a field which met a surge of new contributions over the last years by various authors; see, e.g., Shafer (2021), Vovk and Wang (2021), Ramdas et al. (2023) and Grünwald et al. (2019) amongst others.
Sequential testing methods rely on minimal distributional assumptions and have been used before for sequential forecast evaluation: Henzi and Ziegel (2022) provide e-processes to sequentially assess two probability forecasts for the strong hypothesis that one forecaster is better than the other at all time points. In contrast, Choe and Ramdas (2023) contribute with e-processes for the weaker hypothesis that one forecaster is uniformly better than the other on average. Both studies assess the performance of some forecaster at time $t\in\mathbb{N}$ with respect to the (average) expected loss difference given all information until the previous time step, which goes back to the test of equal conditional predictive ability proposed by Giacomini and White (2006). In contrast to the unconditional approach, most prominently in the seminal work of Diebold and Mariano (1995), the conditional approach uses available information to infer which forecaster is more accurate at a specific date. In this article, we follow the conditional approach and study the strong as well as the weak hypothesis which we further divide into two subhypotheses. In contrast to all aforementioned contributions on forecast comparison, our methods allow to simultaneously draw inference on the performance of $m\geq 2$ different forecasters rather than just being able to compare two different forecasters.
Although the article mainly focuses on forecast evaluation, many of the presented results directly extend to model selection as long as the used information criterion satisfies the imposed assumptions. For this reason, we introduce the problem in a general framework and illustrate the connections to forecast comparison and model selection afterwards. In order to construct sequential model confidence sets, we make use of the e-processes provided by Choe and Ramdas (2023) under the assumption that the observed loss differences are conditionally bounded, which does not only hold for bounded losses but also for prominent choices of scoring rules such as the continuous ranked probability score (CRPS; Matheson and Winkler, 1976) or the quantile score (Gneiting, 2011).
The remainder of the paper is structured as follows. Section 2 presents the problem and introduces sequential model confidence sets. In Section 3, we review important definitions from sequential testing theory and provide methods to construct sequential model confidence sets with respect to three particular notions of superior objects. A simulation study is conducted in Section 4 before we apply the proposed methods in case studies on Covid-19 related deaths and wind gust predictions in Section 5. The main part of the paper closes with a discussion in Section 6. A review on forecast evaluation, technical comments, and additional information for the case studies are available in Appendices A through G.
2 Sequential model confidence sets
We consider a set $\mathcal{M}_{0}=\{1,\dots,m\}$ indexing $m\geq 2$ different statistical models or forecasters, simply referred to as models, and let $L$ be a loss function which measures the quality of the models. We let $L$ be negatively oriented, that is, lower scores correspond to better performance, assume that we collect data in discrete time and denote the loss of model $i\in\mathcal{M}_{0}$ at time $t\in\mathbb{N}$ by $L_{i,t}$. We assume that all random quantities are defined on some underlying measurable space $(\Omega,\mathcal{F})$ equipped with some filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. We denote the family of all probability measures on $(\Omega,\mathcal{F})$ by $\mathfrak{B}(\Omega)$ and write $\mathbb{Q}$ for a generic element in $\mathfrak{B}(\Omega)$.
Example 2.1 (Sequential forecasters).
Let $\{1,\dots,m\}$ denote different forecasters which all sequentially issue predictive distributions or point forecasts $(f_{1,t})_{t\in\mathbb{N}},\dots,(f_{m,t})_{t\in\mathbb{N}}$ for some unknown quantity $(Y_{t})_{t\in\mathbb{N}}\subseteq\mathbb{R}$. For each $i=1,\dots,m$ and $t\in\mathbb{N}$, the forecast $f_{i,t}$ is predictable and refers to the observation $Y_{t}$. To assess the forecasters, one should employ proper scoring rules or consistent scoring functions, see Appendix A. Then, the time series $(L_{i,t})_{t\in\mathbb{N}}$ emerges by applying a proper scoring rule (consistent scoring function) $L$ to the forecast-observation pairs $(f_{i,t},Y_{t})_{t\in\mathbb{N}}$, that is $L_{i,t}=L(f_{i,t},Y_{t})$, $i=1,\dots,m,t\in\mathbb{N}$.
Example 2.2 (Sequential model selection).
Let $\{1,\dots,m\}$ consist of nested and non-nested regression models for some sequentially revealed response $(Y_{t})_{t\in\mathbb{N}}$. At each $t\in\mathbb{N}$, model $i$ observes the covariate(s) $\boldsymbol{X}_{i,t}\subseteq\mathbb{R}^{p_{i}}$, $p_{i}\geq 1,$ and proclaims $Y_{t}=\boldsymbol{\beta}_{i,t}^{\top}\boldsymbol{X}_{i,t}+\varepsilon_{i,t}$ for some $\boldsymbol{\beta}_{i,t}\subseteq\mathbb{R}^{p_{i}}$ and random error $\varepsilon_{i,t}$. The quality of the models is typically evaluated with respect to some likelihood-based information criterion (Ding et al., 2018). For example, for ^$\hat{\ell}_{i,t}$ the empirical log-likelihood of model $i$ given all observations until $t$, $L_{i,t}=-2\hat{\ell}_{i,t}+2p_{i}$ or $L_{i,t}=-2\hat{\ell}_{i,t}+p_{i}\log t$ reveal the well-known Akaike information criterion (AIC) or Bayesian information criterion (BIC), respectively.
Next, we clarify the notion of “best models at time $t\in\mathbb{N}$. For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, consider
$d_{ij,t}=L_{i,t}-L_{j,t},\quad\mu_{ij,t}=\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})$
and their averages
$\hat{\Delta}_{ij,t}=\frac{1}{t}\sum_{s=1}^{t}d_{ij,s},\quad\Delta_{ij,t}=\frac{1}{t}\sum_{s=1}^{t}\mu_{ij,s},$
where we omit the dependence of $\mu_{ij,t}$ and $\Delta_{ij,t}$ on the choice of the probability measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$ for the sake of clarity. Whereas Diebold and Mariano (1995) define model (forecast) superiority with respect to the unconditional expected loss differences, we follow Giacomini and White (2006), Lai et al. (2011), Henzi and Ziegel (2022) and Choe and Ramdas (2023) by defining the superior models in terms of the (average) conditional expected loss differences. In particular, we consider the strongly superior objects
$\mathcal{M}^{\textrm{s},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij,r}\leq 0\textrm{ for all }r\leq t,\textrm{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N},$ (1)
the uniformly weakly superior objects
$\mathcal{M}^{\textrm{uw},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,r}\leq 0\textrm{ for all }j\in\mathcal{M}_{0},\textrm{ for all }r\leq t\right\},\quad t\in\mathbb{N},$ (2)
and the weakly superior objects
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,t}\leq 0\textrm{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$ (3)
For each $t\in\mathbb{N}$, $\mathcal{M}^{\textrm{s},\star}_{t},\mathcal{M}^{\textrm{uw},\star}_{t}$ and $\mathcal{M}^{\textrm{w},\star}_{t}$ are $\mathcal{F}_{t-1}$-measurable random sets with $\mathcal{M}^{\textrm{s},\star}_{t}\subseteq\mathcal{M}^{\textrm{uw},\star}_{t}\subseteq\mathcal{M}^{\textrm{w},\star}_{t}$. Importantly, $\mathcal{M}^{\textrm{w},\star}_{t}\neq\emptyset$ for all $t\in\mathbb{N}$, whereas $\mathcal{M}^{\textrm{s},\star}_{t}$ and $\mathcal{M}^{\textrm{uw},\star}_{t}$ may eventually become empty for all $t\geq t_{0}$ for some $t_{0}\in\mathbb{N}$.
Let $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ be the targeted sequence of superior objects. Of particular interest for this article are the sequences given at (1), (2) and (3), but one could also think of other possible notions of best models. For a given confidence level $\alpha\in(0,1)$, we call (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a sequence of model confidence sets or sequential model confidence sets, for short SMCS(s), for $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}$ at level $\alpha$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}\left(\exists t\geq 1:\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t}\right)\leq\alpha,$ (4)
or equivalently, if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}\left(\forall t\geq 1:\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t}\right)\geq 1-\alpha.$ (5)
We refer to the equivalent conditions (4) and (5) as time-uniform coverage guarantees of the sequential model confidence sets and highlight the fact that the time quantifiers are inside of the probability, which is a much stronger requirement than assuming $\mathbb{Q}(\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t})\geq 1-\alpha$ for all $t\in\mathbb{N}$ (or assuming that there exists a $t\in\mathbb{N}$ with $\mathbb{Q}(\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$); see, e.g., Howard et al. (2021, Lemma 3).
2.1 Discussion of the different sets of superior models
Studying different notions of superior models is important, since, depending on the particular situation, one might interpret the term “best” differently. In this section, we discuss the hypotheses which correspond to the sequences of superior models given at (1), (2), and (3), respectively, in more detail, and give concrete examples for each of them.
If it is reasonable to assume that some models outperform all other models at all time points (in terms of the conditional expected score differences), then we should try to find estimators for $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$. We refer to this assumption as the strong hypothesis. The strong hypothesis particularly applies if we assume independent and identically distributed scores. Then, one typically studies the unconditional expected score differences $\mathbb{E}(d_{ij,t})$ independent of $t\in\mathbb{N}$, as it is done by Diebold and Mariano (1995) or Hansen et al. (2011), who assume stationary performances of the models. Another important instance where the strong hypothesis applies is if the models have nested information sets, as it is e.g. the case in Example 2.2 if the different sets of covariates are nested. Then, the regression models with many covariates typically estimate the given data more accurately, and the AIC or BIC take the corresponding complexities into account by a penalty on the regression dimension. In Section 5.1, we construct SMCSs for different Covid-19 related deaths where the assumption of nested information sets seems plausible.
On the other hand, if we assume that some models have a lower conditional expected score on average rather than at all single time points, then we should target $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ under the uniformly weak hypothesis. Choe and Ramdas (2023, Section 4.4.) argue convincingly that there are many situations where testing for the strong hypothesis may be misleading. Inspired by their arguments, Example 2.4 below considers a forecaster which is ideal on most of the days and only slightly worse than the other forecasters on Sundays. Any powerful method to monitor the strongly superior models over time will exclude this forecaster from the SMCS since she is indeed not strongly superior. Nevertheless, the forecaster is superior under the uniformly weak hypothesis as she is generally better than all other forecasters and only slightly worse than some of them on Sundays.
Finally, we say that the weak hypothesis applies if we expect that also the average performances might change over time. The weak hypothesis is of particular interest if we assume that the models might evolve over time and improve their outputs individually by using past observations. In Section 5.2, we study post-processing methods for wind gust predictions. The results show that in this application it is indeed beneficial to monitor the weakly superior objects as some models which are excluded from the SMCS at earlier time points are included again at later stages, either due to systematic changes in the underlying meteorological model, or due to the different adaptive behaviour of the methods. If we had conducted the study under any of the two stronger hypothesis, we could not have observed that some methods become again competitive towards the end of the test period, since, under these hypotheses, a model can not be included again in the SMCS once it is excluded.
Since the weak hypothesis is the most general, one might wonder, why we do not only consider this particular hypothesis as it is implied by the other two hypotheses anyhow? The answer to this question is given in terms of power: Whereas we impose more assumptions on the data-generating process for the former hypotheses, our methods for the latter are generally less powerful if we apply them to construct SMCSs for the strongly (uniformly weakly) superior models. However, if we construct SMCSs for the strongly (uniformly weakly) superior models even though the corresponding hypotheses do not apply, then we may expect misleading results as argued before.
To conclude, we introduce the simulation settings from Section 4 to give concrete examples for each of the discussed hypotheses.
Example 2.3 (Simulation 1).
For $n=1000$, we sample $(y_{t})_{t=1}^{n}$, where $y_{t}\sim\mathcal{N}(y_{t-1},1),$ for $t=1,\dots,n$ and $y_{0}=0$. We consider $m=49$ different forecasters $\{1,\dots,m\}$ which all sequentially issue predictive distributions for $(y_{t})_{t=1}^{n}$ given by
$f_{i,t}=\mathcal{N}\left(y_{t-1}+\varepsilon_{i},1+\delta_{i}\right),\quad t=1,\dots,n,$
for $(\varepsilon_{i},\delta_{i})\in\{-0.6,-0.4,-0.2,0,0.2,0.4,0.6\}^{2}$. For these forecasters with different biases and dispersion errors, there is exactly one ideal forecaster ${i_{0}}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$. We assess the different forecasters with respect to the continuous ranked probability score (CRPS; Matheson and Winkler (1976)), see Appendix A. Since the CRPS is a proper scoring rule, it follows that $\mathcal{M}^{\textrm{s},\star}_{t}=\{i_{0}\}$ for all $t\in\mathbb{N}$.
Example 2.4 (Simulation 2).
We consider the same data-generating mechanism and the same forecasters as in Example 2.3 with the only exception that the forecaster ${i_{0}}$ now issues the forecasts $f_{i_{0},t}=\mathcal{N}\left(y_{t-1}+\varepsilon_{i_{0},t},1+\delta_{i_{0},t}\right)$, $t=1,\dots,n$, for
$\varepsilon_{i_{0},t}=\delta_{i_{0},t}=\begin{cases}0.3&,t\in 7\mathbb{N}\\ 0&,\textrm{else}.\end{cases}$
That is, forecaster $f_{i_{0}}$ is still ideal on, say, weekdays and Saturdays, however, on Sundays, some forecasters have a smaller bias and dispersion error, and hence $\mathcal{M}^{\textrm{s},\star}_{t}=\emptyset$ for all $t\geq 7$. However, ${i_{0}}$ is uniformly weakly superior, that is $\mathcal{M}^{\textrm{uw},\star}_{t}=\{{i_{0}}\}$ for all $t\in\mathbb{N}$.
Example 2.5 (Simulation 3).
We sample i.i.d. standard normally distributed observations $(y_{t})_{t=1}^{n}$, for $n=800$, and compare $m=3$ different forecasters $\{1,2,3\}$ which issue median predictions $m_{i,t}=y_{t}+\varepsilon_{i,t}$, $i=1,2,3$, for
$\varepsilon_{1,t}=\beta,\quad\varepsilon_{2,t}=\gamma^{t},\quad\varepsilon_{3,t}=\delta t,\quad t=1,\dots,n,$ (6)
for $\beta,\delta>0$ and $10<\gamma<$ 1. We call forecaster $1$ constantly biased, forecaster $2$ improving and forecaster $3$ disimproving. We assess them by the scoring function $L(m,y)=0.5(\Phi(m)-\Phi(y))$, for $m,y\in\mathbb{R}$ and $\Phi$ the cdf of the standard normal distribution, which is a consistent scoring function for the median, see Appendix A. If we choose the parameters as given in Section 4, we have $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$. That is, whereas the disproving forecaster is initially superior, the constantly biased forecaster catches up after some time until the improving forecaster becomes the best in the end.
3 Construction of sequential model confidence sets
In this section, we provide methods to construct SMCSs for the superior models given at (1), (2) and (3). Our constructions build on e-processes, confidence sequences and sequential multiple testing methods as introduced in the following.
3.1 Sequential testing methods and multiple testing
Let $(\Omega,\mathcal{F})$ be a measurable space and $\mathfrak{B}(\Omega)$ be the family of all probability measures on $(\Omega,\mathcal{F})$. A statistical (null) hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a set of probability measures that are potential candidates for the true probability measure $\mathbb{P}$ governing the data generating process. We let $(\mathcal{F}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{F}$ be a filtration and interpret $\mathcal{F}_{t}$ as the available information at time $t\in\mathbb{N}$. We call a process $(M_{t})_{t\in\mathbb{N}}$ a test (super-)martingale for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ if it is a nonnegative (super-)martingale with $M_{0}\leq 1$ for each $\mathbb{Q}\in\mathcal{H}$. Test martingales allow for a financial interpretation in terms of betting as the accumulated wealth of a gambler after having bet a number of times against the null hypothesis (Shafer, 2021). If the null hypothesis is true, then one cannot expect to gain money over time, and thus, large values of the test martingale give evidence against the null hypothesis. More rigorously, any test supermartingale $(M_{t})_{t\in\mathbb{N}}$ for $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ satisfies
$\sup_{\mathbb{Q}\in\mathcal{H}}\mathbb{Q}\left(\exists t\in\mathbb{N}:M_{t}\geq\frac{1}{\alpha}\right)\leq\alpha,\quad\textrm{for all }\alpha\in(0,1).$ (7)
The inequality at (7), referred to as Ville’s inequality (Ville, 1939), follows from the optional stopping theorem for martingales and forms a time-uniform extension of Markov’s inequality. It lays the foundation for safe anytime-valid inference by enabling the construction of time-uniform confidence sequences, sequential tests, or anytime-valid p-values amongst others (Ramdas et al., 2023).
For some problems, it is theoretically impossible to construct nontrivial test martingales (Ramdas et al., 2022). However, often one still can draw sequential inference by using so-called e-processes: An e-process for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a nonnegative adapted stochastic process $(E_{t})_{t\in\mathbb{N}}$ with
$\mathbb{E}_{\mathbb{Q}}(E_{\tau})\leq 1$
for all $\mathbb{Q}\in\mathcal{H}$ and all (possibly infinite) stopping times $\tau$. In other words, for any random stopping time $\tau$, $E_{\tau}$ is an e-variable (Vovk and Wang, 2021). Equivalently, an adapted nonnegative process $(E_{t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}$ if it is upper bounded by a test supermartingale for each $\mathbb{Q}\in\mathcal{H}$ and hence, any test supermartingale is also an e-process and Ville’s inequality continues to hold for e-processes (Ramdas et al., 2020).
Test martingales and e-processes allow for sequential testing. A sequential testing procedure or sequential test for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is an adapted stochastic process $(\Psi_{t})_{t\in\mathbb{N}}\subseteq\{0,1\}$, where $\Psi_{t}=1$ indicates that the hypothesis $\mathcal{H}$ is rejected at $t\in\mathbb{N}$. A sequential test $(\Psi_{t})_{t\in\mathbb{N}}$ for $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is said to uniformly control the type-I error at level $\alpha\in(0,1)$ if
$\sup_{\mathbb{Q}\in\mathcal{H}}\mathbb{Q}(\exists t\in\mathbb{N}:\Psi_{t}=1)\leq\alpha.$
By Ville’s inequality at (7), any e-process $(E_{t})_{t\in\mathbb{N}}$ directly leads to a sequential testing procedure $(\Psi_{t})_{t\in\mathbb{N}}$ with uniform type-I error control at level $\alpha\in(0,1)$ by defining
$\Psi_{t}=\mathbbm{1}\{E_{t}\geq 1/\alpha\},\quad t\in\mathbb{N}.$ (8)
If we test more than one hypothesis simultaneously, then we have to correct for multiple testing. In the spirit of Vovk and Wang (2021), we call $m\geq 2$ e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ family-wise valid (FWV) if there exists a dominating family of e-processes, that is, if there exists a family $\{(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ is an e-process for ${}\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and for any $i=1,\dots,m$ with $\mathbb{Q}\in\mathcal{H}_{i}$, $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ upper bounds $(E_{i,t})_{t\in\mathbb{N}}$. In Appendix C, we show that for family-wise valid e-processes, the corresponding sequential tests given at (8) time-uniformly control the family-wise error rate (FWER).
To derive SMCS for the sequence of weakly superior objects, we will make use of confidence sequences with time-uniform coverage guarantees (Howard et al., 2021; Ramdas et al., 2023). For $\alpha\in(0,1)$, a (time-uniform) $(1-\alpha)$-confidence sequence for a stochastic process $(\theta_{t})_{t\in\mathbb{N}}\subseteq\Theta$ in some space $\Theta$ is a sequence $(C_{t})_{t\in\mathbb{N}}\subseteq 2^{\Theta}$ such that
$\mathbb{Q}\left(\forall t\in\mathbb{N}:\theta_{t}\in C_{t}\right)\geq 1-\alpha,\quad\textrm{for all }\mathbb{Q}\in\mathfrak{B}(\Omega).$
3.2 SMCSs for the strongly superior models
In this section, we construct SMCSs for the sequence of strongly superior models $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$ given at (1) by a sequential testing procedure. Throughout this section, we assume that for any $i,j\in\mathcal{M}_{0}$, the sequence of loss differences $(d_{ij,t})_{t\in\mathbb{N}}$ is conditionally bounded with respect to some predictable sequence $(c_{ij,t})_{t\in\mathbb{N}}\subseteq\mathbb{R}_{+}$, that is $|d_{ij,t}|\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$. For $i,j\in\mathcal{M}$, we consider the strong hypotheses
$\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,t}\leq 0,\forall t\in\mathbb{N}\}$
and the strong intersection hypotheses
$\mathcal{H}^{\textrm{s}}_{i\cdot}=\bigcap_{j\in\mathcal{M}_{0}}\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,t}\leq 0,\forall t\in\mathbb{N},\forall j\in\mathcal{M}_{0}\}.$
Then, for any predictable $(\lambda_{ij,t})_{t\in\mathbb{N}}$ with $0\leq\lambda_{ij,t}\leq 1/c_{ij,t}$, we have that
$M_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}(d_{ij,r}-\mu_{ij,r})),\quad t\in\mathbb{N}$
is a nonnegative martingale. Importantly, for any $\mathbb{Q}\in\mathcal{H}^{\textrm{s}}_{ij}$, $M_{ij,t}$ upper bounds
$E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r}),\quad t\in\mathbb{N},$
Hence, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}^{\textrm{s}}_{ij}$, and thus also for $\mathcal{H}^{\textrm{s}}_{i\cdot}$. For $i=1,\dots,m$, we let
$E_{i\cdot,t}=\frac{1}{m-1}\sum_{j\neq i}E_{ij,t},\quad t\in\mathbb{N},$
which is also an e-process for $\mathcal{H}^{\textrm{s}}_{i\cdot}$ as the arithmetic mean of $m-1$ e-processes. Finally, we adjust the e-processes
$(E_{1\cdot t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}},$ (9)
symmetrically by the closure principle in the spirit of Vovk and Wang (2021), as explained in detail in Appendix C, to obtain family-wise valid e-processes $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}^{\textrm{s}}_{1\cdot},\dots,\mathcal{H}^{\textrm{s}}_{m\cdot}$. We use the arithmetic mean as the $e$-merging function as it essentially dominates all symmetric e-merging functions (Vovk and Wang, 2021, Proposition 3.1.). For $\alpha\in(0,1)$, we apply the sequential tests given at (8) and let
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}^{\star}\leq 1/\alpha\right\}.$ (10)
That is, at each $t\in\mathbb{N}$, we exclude model $i\in\mathcal{M}_{0}$ from the SMCS if we may reject the hypothesis $\mathcal{H}^{\textrm{s}}_{i\cdot}$ according to the sequential testing procedure, and include it otherwise.
Lemma 3.1.
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (10) is a sequence of model confidence sets at level $\alpha$ for the sequence of strongly superior objects $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$, and so is its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$.
Proof.
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{s},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i\cdot,t}^{\star}>1/\alpha\textrm{ for some }i\in\mathcal{M}^{\textrm{s},\star}_{t}\right).$
Moreover, $i\in\mathcal{M}^{\textrm{s},\star}_{t}$ if and only if $\mathbb{Q}\in\mathcal{H}^{\textrm{s}}_{i\cdot,t}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,r}\leq 0,\forall r\leq t,\forall j\in\mathcal{M}_{0}\}$. Since $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ are FWV, there exists a dominating family of e-processes $\{(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $E_{r}^{\mathbb{Q}}\geq E_{i\cdot,r}^{\star}$ for all $r\leq t$ and all $i\in\mathcal{M}^{\textrm{s},\star}_{t}$. So, we may conclude
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{s},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})$ $=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i\cdot,t}^{\star}>1/\alpha\textrm{ for some }i\in\mathcal{M}^{\textrm{s},\star}_{t}\right)$
$\leq\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{t}^{\mathbb{Q}}>1/\alpha\right)\leq\alpha.$
It follows directly, that the running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$, is a SMCS as well, since $i\notin\mathcal{M}^{\star}_{t_{0}}$ implies $i\notin\mathcal{M}^{\star}_{t}$ for all $t\geq t_{0}.$
Remark 1.
Alternatively, one could adjust the $m\cdot(m-1)$ e-processes $(E_{ij,t})_{t\in\mathbb{N}}$ directly to obtain FWV e-processes $(E^{\star}_{ij,t})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}^{\textrm{s}}_{ij}$ and exclude model $i$ from the SMCS at $t\in\mathbb{N}$ if $E^{\star}_{ij,t}>1/\alpha$ for some $j\neq i$. However, the corresponding tests are less powerful, since for any $i\in\mathcal{M}_{0}$ and $j\neq i$,
$E^{\star}_{ij,t}$ $=\min\left\{\frac{1}{\lvert I\rvert}\sum_{(k,l)\in I}E_{kl,t}\;\Big{|}\;I\subseteq\{1,\dots,m\}^{2},(i,j)\in I\right\}$
$\leq\min\left\{\frac{1}{\lvert I\rvert(m-1)}\sum_{k\in I}\sum_{l=1}^{m}E_{kl,t}\;\Big{|}\;I\subseteq\{1,\dots,m\},i\in I\right\}=E_{i\cdot}^{\star}.$
Remark 2.
Apart from the FWER, the false discovery rate (FDR) is arguably the most commonly used criterion in multiple testing (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001). Recently, Wang and Ramdas (2022) studied FDR control for e-values, which turn out to be particularly suitable for this problem due to their behaviour under taking expectations. In Appendix D, we demonstrate how one could alternatively construct SMCSs with a time-uniform control of the FDR, which is generally less restrictive than FWER.
3.3 SMCSs for the uniformly weakly superior objects
We construct SMCSs for the sequence of uniformly weakly superior objects $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ defined at (2) analogously as in the previous section. For simplicity, we assume first that for any $i,j\in\mathcal{M}$, there exists a constant $c_{ij}>0$ with $\lvert d_{ij,t}\rvert\leq c_{ij}/2$ for all $t\in\mathbb{N}$. In Section 3.5, we explain how we can weaken this assumption to conditionally bounded loss differences.
For $i,j\in\mathcal{M}$, consider the weak hypotheses
$\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,t}\leq 0,\forall t\in\mathbb{N}\}$
and the weak intersection hypotheses
$\mathcal{H}^{\textrm{uw}}_{i\cdot}=\bigcap_{j\in\mathcal{M}_{0}}\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,t}\leq 0,\forall t\in\mathbb{N},\forall j\in\mathcal{M}_{0}\}.$
Choe and Ramdas (2023) show that, for any $0\leq\lambda_{ij}<1/c_{ij}$,
$M_{ij,t}=\exp\left\{\lambda_{ij}t(\hat{\Delta}_{ij,t}-\Delta_{ij,t})-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad t\in\mathbb{N}$
is a test supermartingale under any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, where
$\psi_{E,c}(\lambda)=(-\log(1-c\lambda)-c\lambda)/c^{2}$ (11)
for $c>0$ and $0\leq\lambda<1/c$ and
$V_{ij,t}=\sum_{r=1}^{t}(d_{ij,r}-\gamma_{ij,r})^{2}$ (12)
for some predictable sequence $(\gamma_{ij,t})_{t\in\mathbb{N}}$ with $\lvert\gamma_{ij,t}\lvert\leq c_{ij}/2$. Importantly, for any $\mathbb{Q}\in\mathcal{H}^{\textrm{uw}}_{ij}$, $M_{ij,t}$ upper bounds
$E_{ij,t}=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij,t}-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\}$ (13)
for all $t\in\mathbb{N}$. That is, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}^{\textrm{uw}}_{ij}$ and hence also for $\mathcal{H}^{\textrm{uw}}_{i\cdot}$. Analogously as before, we adjust the averages $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$, $i=1,\dots,m$, symmetrically by the closure principle to construct SMCSs for $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ by the FWV e-processes $(E_{1\cdot,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t}^{\star})_{t\in\mathbb{N}}$ as in (10), compare Appendix C.
3.4 SMCSs for the weakly superior objects
Our SMCS construction for the weakly superior objects defined at (3) builds on time-uniform confidence regions. For any $i,j\in\mathcal{M}_{0}$, we again assume that there exists $c_{ij}>0$ with $\lvert d_{ij,t}\rvert\leq c_{ij}/2$ for all $t\in\mathbb{N}$ and discuss a possible weakening of this assumption in Section 3.5. For $t\in\mathbb{N}$ and $0\leq\lambda_{ij}\leq 1/c_{ij}$, let
$M_{ij,t}(x)=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij}-\lambda_{ij}tx-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\},\quad x\in\mathbb{R},t\in\mathbb{N},$
where $\psi_{E,c_{ij}}$ and $V_{ij,t}$ are given in (11) and (12), respectively. The following lemma shows the one-to-one correspondence between confidence sequences and sequential tests, and is found e.g. in Howard et al. (2021, Lemma 1) or Waudby-Smith and Ramdas (2024, Theorem 1).
Lemma 3.2.
For any $i,j\in\mathcal{M}_{0}$ and $\alpha\in(0,1)$, the sequence
$C_{t}=\left\{x\in[-c_{ij}/2,c_{ij}/2]\mid M_{ij,t}(x)\leq 1/\alpha\right\},\quad t\in\mathbb{N},$
is a $(1-\alpha)$-confidence sequence for $(\Delta_{ij,t})_{t\in\mathbb{N}}$.
The claim follows directly by Ville’s inequality observing that $(M_{ij,t}(\Delta_{ij,t}))_{t\in\mathbb{N}}$ is a nonnegative supermartingale for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. Note that $M_{ij,t}(x)\leq 1/\alpha$ if and only if $x\geq\hat{\Delta}_{ij,t}-(\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}+\log(1/\alpha))/(\lambda_{ij}t)$. That is, Lemma 3.2 actually yields time-uniform lower confidence bounds for $(\Delta_{ij,t})_{t\in\mathbb{N}}$. Next, we show how to combine these lower confidence bounds simultaneously for all models $i\in\mathcal{M}_{0}$, and how we can use them to construct SMCSs for $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$.
We let $\mathbb{R}_{0}^{m\times m}$ be the family of all $m\times m$-matrices with diagonal entries equal zero and let
$\boldsymbol{c}=\begin{pmatrix}0&c_{12}&\cdots&c_{1m}\\ c_{21}&0&\cdots&c_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ c_{m1}&c_{m2}&\dots&0\\ \end{pmatrix}\in\mathbb{R}_{0}^{m\times m},\quad\boldsymbol{\Delta}_{t}=\begin{pmatrix}0&\Delta_{12,t}&\cdots&\Delta_{1m,t}\\ \Delta_{21,t}&0&\cdots&\Delta_{2m,t}\\ \vdots&\vdots&\ddots&\vdots\\ \Delta_{m1,t}&\Delta_{m2,t}&\cdots&0\\ \end{pmatrix}\in\mathbb{R}_{0}^{m\times m},\quad t\in\mathbb{N}.$
For $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C}\in\mathbb{R}_{0}^{m\times m}$ and $C\in\mathbb{R}$, we understand $\boldsymbol{A}\leq\boldsymbol{B}$ and $C\cdot\boldsymbol{A}$ componentwise, and write $\boldsymbol{B}\leq\boldsymbol{A}\leq\boldsymbol{C}$ equivalently as $\boldsymbol{A}\in[\boldsymbol{B},\boldsymbol{C}]$. For $\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}$, we let
$M_{t}(\boldsymbol{X})=\frac{1}{m(m-1)}\sum_{\begin{subarray}{c}i,j=1,\dots,m\\ i\neq j\end{subarray}}M_{ij,t}(X_{ij}),\quad t\in\mathbb{N}.$
Lemma 3.3.
For any $\alpha\in(0,1)$, the sequence
$C_{t,1-\alpha}=\left\{\boldsymbol{X}\in[-\boldsymbol{c}/2,\boldsymbol{c}/2]\mid M_{t}(\boldsymbol{X})\leq 1/\alpha\right\}$
is a time-uniform $(1-\alpha)$-confidence sequence for the sequence $(\boldsymbol{\Delta}_{t})_{t\in\mathbb{N}}$.
Proof.
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t,1-\alpha})$ $=\mathbb{Q}(\exists t\in\mathbb{N}:M_{t}(\boldsymbol{\Delta}_{t})>1/\alpha)\leq\alpha,$
by Ville’s inequality since $(M_{t}(\boldsymbol{\Delta}_{t}))_{t\in\mathbb{N}}$ is a test supermartingale under any $\mathbb{Q}\in\mathfrak{B}(\Omega)$. ∎
Next, we define
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t,1-\alpha}\cap\left(\mathbb{R}_{0}^{m\times m}\right)_{ij,-}\neq\emptyset\textrm{ for all }j\neq i\right\},\quad t\in\mathbb{N},$ (14)
where $\left(\mathbb{R}_{0}^{m\times m}\right)_{ij,-}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid X_{ij}\leq 0\}$. That is, for each model $i\in\mathcal{M}_{0}$, we check whether $\Delta_{ij,t}$ may potentially be negative for all $j\neq i$ according to the confidence region $C_{t,1-\alpha}$. If there exists $j\neq i$ such that we may reject that $\Delta_{ij,t}\leq 0$ with high probability, then we exclude $i$ from the SMCS.
Lemma 3.4.
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (14) is a a sequence of model confidence sets at level $\alpha$ for the sequence of weakly superior objects $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\in\mathbb{N}}$.
Proof.
By Lemma 3.3, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{w},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})$
$=$ $\mathbb{Q}(\exists t\in\mathbb{N}:i\in\mathcal{M}^{\textrm{w},\star}_{t}\textrm{ and }i\not\in\widehat{\mathcal{M}}_{t}\textrm{ for some }i\in\mathcal{M}_{0})$
$=$ $\mathbb{Q}(\exists t\in\mathbb{N},\exists i\in\mathcal{M}_{0}:\Delta_{ij,t}\leq 0\textrm{ for all }j\neq i\textrm{ and }C_{t}\cap\left(\mathbb{R}^{m\times m}\right)_{i\ell,-}=\emptyset\textrm{ for some }\ell\neq i)$
$\leq$ $\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t})\leq\alpha.$
The following Lemma, which is proved in Appendix E, characterizes the shape of the confidence regions $(C_{t,1-\alpha})_{t\in\mathbb{N}}$ and may be used to compute the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ more efficiently.
Lemma 3.5.
For any $\alpha\in(0,1)$ and $t\in\mathbb{N}$, $C_{t,1-\alpha}$ is a convex upper set. That is, for $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$, it follows that $\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\in C_{t,1-\alpha}$ for all $\lambda\in[0,1]$, and $\boldsymbol{A}\in C_{t,1-\alpha}$ and $\boldsymbol{A}\leq\boldsymbol{C}$ implies that $\boldsymbol{C}\in C_{t,1-\alpha}$ for any $\boldsymbol{C}\leq\boldsymbol{c}/2$.
By Lemma 3.5, we may express ^$\widehat{\mathcal{M}}_{t}$ equivalently as
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid C_{t,1-\alpha}\cap[\boldsymbol{0},\boldsymbol{c}/2]_{ij,0}\neq\emptyset\textrm{ for all }j\neq i\right\},\quad t\in\mathbb{N},$
where $\boldsymbol{0}=(0)_{i,j=1}^{m}\in\mathbb{R}_{0}^{m\times m}$ and $[\boldsymbol{0},\boldsymbol{c}/2]_{ij,0}=\{\boldsymbol{X}\in[\boldsymbol{0},\boldsymbol{c}/2]\mid X_{ij}=0\}$. That is, we exclude model $i\in\mathcal{M}_{0}$ from the SMCS at $t\in\mathbb{N}$, if there exists $j\neq i$ such that the confidence region $C_{t,1-\alpha}$ has an empty intersection with the hyperplane $\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid X_{ij}=0\}$.
3.5 Predictable bounds and betting schemes
The methods provided in Sections 3.3 and 3.4 require that the score differences are uniformly bounded over time, which is clearly more restrictive than the assumption of conditionally bounded score differences imposed in Section 3.2. However, one can always transform conditionally bounded score differences into uniformly bounded differences thereby modifying the underlying loss and the corresponding superior objects.
Assume that $\lvert d_{ij,t}\rvert\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$ for some predictable $(c_{ij,t})_{t\in\mathbb{N}}$. Then, the transformed loss differences $\tilde{d}_{ij,t}=d_{ij,t}/c_{ij,t},t\in\mathbb{N}$ are uniformly bounded with constant $c=1$ that is $\lvert\tilde{d}_{ij,t}\rvert\leq 1/2$ for all $t\in\mathbb{N}$. However, if we use the transformed loss differences for the methods of the previous sections, then we target (possibly) different sequences of superior objects
~$\tilde{\mathcal{M}}^{\textrm{uw},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,r}\leq 0,\forall r\leq t,\forall j\in\mathcal{M}_{0}\right\},\quad\tilde{\mathcal{M}}^{\textrm{w},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\tilde{\Delta}_{ij,t}\leq 0,\forall j\in\mathcal{M}_{0}\right\},$
for
$\tilde{\Delta}_{ij,t}=\frac{1}{t}\sum_{r=1}^{t}\tilde{\mu}_{ij,r}=\frac{1}{t}\sum_{r=1}^{t}\mathbb{E}(\tilde{d}_{ij,r}\mid\mathcal{F}_{r-1})=\frac{1}{t}\sum_{r=1}^{t}\frac{\mu_{ij,r}}{c_{ij,r}}.$
In Section 5, we use the given transformation to convert conditionally bounded CRPS differences of wind gust forecasts into uniformly bounded loss differences. Importantly, for any proper scoring rule (consistent scoring function) ${\rm S}$, the scaled function $\tilde{{\rm S}}={\rm S}/c$ is proper (consistent) as well for any $c>0$, an observation which justifies the transformation from a theoretical perspective. Additionally, with the given scaling, all observations have the same maximal impact on the ranking, an effect which seems appealing for many applications.
Remark 3.
We assume conditionally bounded loss differences in order to work with the empirical Bernstein e-processes at (13) provided by Choe and Ramdas (2023). Clearly, for many important loss functions this assumption is not fulfilled; see, e.g., Example 2.2 where the log-likelihood differences are typically unbounded. However, if the loss differences emerge from (or may be bounded in the tails by) a parametric family of distributions, as it is for example the case in Example 2.2, there are other well-studied e-processes available; see, e.g., Howard et al. (2020), Pérez-Ortiz et al. (2023) or Hao et al. (2024) amongst others. In such situations, our methods directly extend with a suitable choice of e-processes at (13).
We conclude with a comment on the choice of the parameter $\lambda$. For the uniformly weak hypothesis and the weak hypothesis, we assume a universal bound $c>0$ on the loss differences, and $\lambda$ may be any fixed value in the interval $(0,c^{-1})$. If there is no good reason to do otherwise, we suggest to choose $\lambda=(2c)^{-1}$ by default. In betting language, this corresponds to bet half of the accumulated evidence at each time step (Shafer, 2021). For the strong hypothesis, we should choose $\lambda$ predictably and in a smart way in order to increase power. Clearly, there are many possible betting schemes. In Section 5.1, we propose a particular betting scheme which makes use of all previous observations and the assumption that the relative performance of two models does not change too quickly over time. Another option would be the method of mixtures (Robbins, 1970), where we integrate over all possible values of $\lambda$ over a particular probability distribution instead of choosing one specific value. Mixtures of e-processes are again e-processes and may be expressed in closed form for some distributions (Howard et al., 2021, Appendix A.3). The method of mixtures is one of the most widely-studied techniques for constructing uniform bounds which are shown to be unimprovable under certain conditions (Robbins and Siegmund, 1970, Howard et al., 2021). However, often it is analytically opaque and computationally tedious, and does not lead to closed-form confidence sequences. In accordance with Waudby-Smith and Ramdas (2024), for the strong hypothesis, our experience shows that choosing $\lambda$ predictably in a smart ways is generally more promising than the methods of mixtures. Also for the weaker hypotheses, we could not improve the power of our methods significantly by using mixtures instead of choosing a fixed $\lambda$ value. For this reason, and to make the computations more tractable, we suggest to work with the given e-processes directly rather than with the more complicated mixture e-processes also for both weak hypotheses.
4 Simulations
4.1 Simulation 1
We consider the setup given in Example 2.3. That is, we consider $n=1000$ realizations from a sequence $Y_{t}\sim\mathcal{N}(Y_{t-1},1)$ with $Y_{0}=0$, and assess $m=49$ forecasters which issue sequentially probabilistic predictions with respect to the CRPS. We write $Y_{t}=Y_{t-1}+Z_{t}$, for an i.i.d.  standard normally distributed sequence $(Z_{t})_{t\in\mathbb{N}}$, and use the formulas provided by Grimit et al. (2006) to derive
$L_{i,t}$ $=\textrm{CRPS}(f_{i,t},Y_{t})$
$=\sqrt{1+\delta_{i}}\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\left(2\Phi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-1\right)+2\varphi\left(\frac{Z_{t}-\varepsilon_{i}}{\sqrt{1+\delta_{i}}}\right)-\frac{1}{\sqrt{\pi}}\right),$ (15)
where $\Phi$ and $\varphi$ denote the cdf and the density function of the standard normal distribution, respectively. Let $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be the canonical filtration, that is $\mathcal{F}_{t}$ is generated by all observations $Y_{r}$ and forecasts $f_{i,r}$, $i=1,\dots,m$, $r\leq t$. For any $t\in\mathbb{N}$, we have $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mathbb{E}(d_{ij,1})$ by independence. Since the CRPS is a proper scoring rule, we conclude
$\mathcal{M}^{\textrm{s},\star}_{t}=\mathcal{M}^{\star}=\left\{i\in\mathcal{M}_{0}\mid\mathbb{E}(d_{ij,1})\leq 0,\forall j\in\mathcal{M}_{0}\right\}=\{i_{0}\}$
for the forecaster $i_{0}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$.
We construct SMCSs for $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\leq n}$ as proposed in Section 3.2 with confidence level $\alpha=0.1$. We let $\lambda_{ij,t}=(2c_{ij})^{-1}$ for the universal bound $c_{ij}>0$ on the score differences $(d_{ij,t})$ given in Appendix B. Figure 1 shows the average size of the SMCS over time for $N=100$ realizations of the simulation. Importantly, we have a 100% coverage rate of the best model $i_{0}$, that is, in all runs of the simulation, the SMCS never wrongly excludes $i_{0}$.
Figure 1: The average number of models in the SMCS in simulations 1 and 2 over $N=100$ repetitions. In simulation 1, the size of the SMCS tends to the actual number of superior objects as the sample size increases, whereas there are still approximately $10$ models remaining in the SMCS for the uniformly weakly superior objects after $1000$ observations. For both simulations, the SMCS never wrongly excludes the best model $i_{0}$.
4.2 Simulation 2
We consider the simulation setting as given in Example 2.4. By (4.1),
$\mathbb{E}(L_{i,t})$ $=\mathbb{E}\big{(}Z_{t}\Phi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}-\varepsilon_{i}\mathbb{E}\big{(}\Phi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}+\varepsilon_{i}+2\sqrt{1+\delta_{i}}\mathbb{E}\big{(}\varphi_{\varepsilon_{i},1+\delta_{i}}(Z_{t})\big{)}-\sqrt{\frac{1+\delta_{i}}{\pi}}$
$=\frac{1+2\sqrt{1+\delta_{i}}}{\sqrt{2\pi(2+\delta_{i})}}\exp\left\{-\frac{\alpha_{i}^{2}}{2}\right\}-\varepsilon_{i}\Phi\left(-\alpha_{i}\right)+\varepsilon_{i}-\sqrt{\frac{1+\delta_{i}}{\pi}},$ (16)
where $\alpha_{i}=\varepsilon_{i}/\sqrt{2+\delta_{i}}$ and $\Phi_{\mu,\sigma^{2}}$ ($\varphi_{\mu,\sigma^{2}}$) denotes the cdf (density) of $\mathcal{N}(\mu,\sigma^{2})$ for $\mu\in\mathbb{R}$ and $\sigma>0$. For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, let $\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})=\mathbb{E}(d_{ij,t})=\mu_{ij,t}$. Since the CRPS is a proper scoring rule, we have $\mu_{i_{0}j,t}\leq 0$ for all $t\notin 7\mathbb{N}$, for all $j\in\mathcal{M}_{0}$. By (16), for $j\in\mathcal{M}_{0}$ with $(\varepsilon_{j},\delta_{j})\in\{-0.2,0.2\}^{2}$, we have $\mu_{i_{0}j,t}>0$ for all $t\in 7\mathbb{N}$. Thus, $\mathcal{M}^{\textrm{s},\star}_{t}=\emptyset$ for $t\geq 7$. However, equation (16) allows to show numerically that $\Delta_{i_{0}j,t}\leq 0$ for all $j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, and thus $\mathcal{M}^{\textrm{uw},\star}_{t}=\{i_{0}\}$ for all $t\in\mathbb{N}$.
We compute SMCSs for $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\leq n}$ as suggested in Section 3.3 with confidence level $\alpha=0.1$. We transform the conditionally bounded CRPS differences (see Appendix B) as proposed in Section 3.5 into uniformly bounded score differences $\lvert\tilde{d}_{ij,t}\rvert\leq C/2$ for $C=2$, and let $\lambda=(2C)^{-1}=1/4$. Figure 1 displays the average size of the SMCS for $N=100$ repetitions of the simulation. Again, we have a 100% coverage rate of the superior model $i_{0}$. Notably, we have less power compared to the SMCS for the strongly superior objects and still remain with approximately $10$ surviving models after $1000$ observations.
4.3 Simulation 3
We consider the simulation setting as given in Example 2.5. Let $i=1,2,3$ be forecasters issuing median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, $t\in\mathbb{N}$, for i.i.d. standard normally distributed $(Y_{t})_{t\in\mathbb{N}}$, where the biases $\varepsilon_{i,t}$ are given in (6). For $t\in\mathbb{N}$, we have by independence
$2\mu_{ij,t}=\mathbb{E}(\Phi(m_{i,t}))-\mathbb{E}(\Phi(m_{j,t}))=\Phi\left(\frac{\varepsilon_{i,t}}{\sqrt{2}}\right)-\Phi\left(\frac{\varepsilon_{j,t}}{\sqrt{2}}\right),\quad i,j=1,2,3,$
and thus
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i=1,2,3\mid\sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{i,t}}{\sqrt{2}}\right)\leq\sum_{r=1}^{t}\Phi\left(\frac{\varepsilon_{j,t}}{\sqrt{2}}\right),j=1,2,3\right\}.$ (17)
We let $\beta=0.6,\gamma=0.998$ and $\delta=0.008$. Then, one can numerically show by (17) that $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$.
We sample $n=800$ observations and construct SMCSs for $(\mathcal{M}^{\textrm{w},\star}_{t})_{t\leq n}$ at level $\alpha=0.1$ as suggested in Section 3.4. We use the fact that $(d_{ij,t})_{t\in\mathbb{N}}$ is uniformly bounded with $c=1$ and let $\lambda_{ij}=1.1^{-1}$, $i,j=1,2,3$. Figure 2 shows the accumulated observed losses and the resulting SMCS for one realization of the simulation (left), and the average size of the SMCS over time for $N=100$ runs (right). Again, the SMCS includes the theoretical superior objects given at (17) at all time points.
Refer to captionRefer to caption
Figure 2: For $i=1,2,3$, the left panel shows the realized accumulated losses $\sum_{r=1}^{t}L_{i,r}$, $t\leq n$, for one simulation. The black vertical lines indicate $t=154$ and $t=550$. The resulting SMCS is given in the upper part with the respective colors: The SMCS correctly excludes the improving forecaster (blue) at the beginning and the disimproving forecaster (green) relatively quickly after the first change point until it includes the improving forecaster again and finally excludes the constantly biased forecaster (red). The right panel shows the average size of the SMCS over $N=100$ realizations, where we again have a $100\%$ coverage rate of the superior models.
5 Case Studies
5.1 Covid-19 case study
After the outbreak of the Covid-19 pandemic, the Reich lab at the University of Massachusetts collaborated with the United States centers for disease control and prevention (CDC) to create the Covid-19 forecast hub, a repository containing point and probabilistic forecasts for incident cases, hospitalizations, and Covid-19 related deaths. The repository was founded in March 2020, and attracted the submission of forecasts by more than $60$ different teams. The collection of forecasts is still ongoing, although most of the models had stopped submitting forecasts by January 2024, due to the subsidence of the pandemic.
5.1.1 The data
The dataset, which we use in this application, is discussed in detail in Cramer et al. (2022a) and Cramer et al. (2022b), and publicly available at the Covid-19 forecast hub GitHub repository and on the Zoltar forecast archive. It contains forecasts with corresponding observations for key epidemiology indices for 55 different locations in the US as well as on the aggregated US national level. The forecasts are issued by a total of 69 different forecast models and reported at $23$ different quantile levels with forecast horizons ranging from $1$ day to $20$ weeks. Importantly, many models are not directly comparable since not all of them have issued their forecasts on the same weekdays or throughout the whole period.
In this case study, we consider 1-week-ahead forecasts for Covid related deaths on the national level, issued weekly in the period from 06/06/2020 to 04/03/2024, and focus on the comparison of the following $m=6$ different forecasting models: Firstly, we consider the benchmark model baseline, which naively issues the most recent outcome as the median prediction for the following week and forms a predictive distribution around this median prediction based on the past weekly incidences. Secondly, we consider two epidemic models PSI-draft and MOBS-gleam issuing their forecasts based on epidemiological assumptions, and the neural network based model GT-deep, which was the first purely data-driven model to be included in the Covid-19 forecast hub. Finally, we consider the summary models ensemble and CDC-ensemble, where the former combines the most recent predictions of all other models into one predictive distribution, and the latter is supposed to improve the former by only considering the 10 current best models, measured by the average weighted interval score over the 12 most recent weeks. For a more detailed documentation of the models, we refer to Appendix F.
Regarding the nature of the models, we perform our comparison under the strong hypothesis, since we expect the ensemble models to have consistent superior performance over the individual models, which in turn are expected to be better than the naive baseline model issuing its forecasts only based on the previous outcomes and on no additional information. Furthermore, we can argue that in such a highly sensible prediction task, we should seek to detect the models that issue accurate predictions consistently over time and not just on average.
Notably, the outcomes as well as the forecasts in the given dataset have a very unstable behaviour and exhibit large variation in scale, see e.g. the upper left panel of Figure 3, which suggests to apply a log transformation on all quantities of interest. Moreover, the forecast differences are typically not very distinctive for the median, whereas they become more apparent when we look at the tails, compare to the lower panels in Figure 3.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Time progression of predicted and actual Covid related mortality in linear (upper left) and log scale (upper right). For the forecasts at the tails, the differences between the different models are more pronounced than for the median, see the lower panels for $\tau=0.15,0.975$.
5.1.2 Methods and implementation details
For $\tau\in(0,1)$, we assess the different quantile forecasts at level $\tau$ with respect to the generalized piecewise-linear (GPL) quantile scoring function (Gneiting, 2011),
$S_{\tau}(x,y)=\left(\mathbf{1}\{x\geq y\}-\tau\right)\left(\log(x)-\log(y)\right),\quad x,y\in\mathbb{R}.$
The log transformation may thus easily be justified from a statistical point of view by simply setting $g=\log$ in the formula of the GPL score, see Appendix A for details. In Appendix B, it is shown that the resulting quantile score differences are conditionally bounded. More precisely, for $\tau$-quantile predictions $(x_{i,t})_{t\in\mathbb{N}},(x_{j,t})_{t\in\mathbb{N}}\subseteq\mathbb{R}$ issued by the models $i,j=1,\dots,m$, we have $\lvert d_{ij,t}\rvert=\lvert S_{\tau}(x_{i,t},y_{t})-S_{\tau}(x_{j,t},y_{t})\rvert\leq c_{ij,t}/2$ for
$c_{ij,t}=2\max\left\{\tau,1-\tau\right\}\lvert\log(x_{1,t})-\log(x_{2,t})\rvert,\quad t\in\mathbb{N}.$ (18)
We apply our methods of Section 3.2 to construct SMCSs for the strongly superior objects, where we use the e-processes $E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r})$ for $(\lambda_{ij,t})_{t\in\mathbb{N}}$ predictable with $0<\lambda_{ij,t}<c_{ij,t}^{-1}$, $t\in\mathbb{N}$. For $i,j=1,\dots,m$ and $t\in\mathbb{N}$, we suggest to set
$\lambda_{ij,t}=\frac{1}{K_{ij,t}c_{ij,t}+\varepsilon}$
for some small $\varepsilon>0$, included as a safeguard for the scenario where the two forecaster issue the same prediction and thus $c_{ij,t}=0$, and for
$K_{ij,t}=\frac{2-\left|\tau-\frac{1}{2}\right|}{1+\left|\tau-\frac{1}{2}\right|}\cdot\frac{3\pi/2+\arctan(-d_{ij,t-1})}{\pi}\geq 1.$
The coefficient $K_{ij,t}$ depends on the sign and magnitude of the most recent score difference $d_{ij,t-1}$ as well as of the centrality of $\tau$: If $d_{ij,t-1}>0$, then we have evidence against the null hypothesis that model $i$ is strictly superior than model $j$, and the quantity $K_{ij,t}$ becomes small. That is, the parameter $\lambda_{ij,t}$ tends to be larger, and we bet more against the null hypothesis. On the other hand, $K_{ij,t}$ becomes larger for $d_{ij,t-1}<0$, which leads to smaller values of the tuning parameters $\lambda_{ij,t}$. Finally, the primary multiplicative factor, which takes values in $[1,2]$, is used to reduce the influence of the term $\max\left\{\tau,1-\tau\right\}$ on the bound in (18). It converges to $1$ as $\tau\downarrow 0$ or $\tau\uparrow 1$, and to $2$ as $\tau\to 0.5$. This means that we bet more at the tails where the forecasts are typically more distinctive and less if $\tau$ is close to $0.5$ where the forecasts are typically less distinguishable.
Choosing a smart betting strategy has an effect on the power of the methods as discussed in Section 3.5. In Appendix F, we compare the suggested choice of the predictable $\lambda_{ij,t}$’s with the naive approach where $K_{ij,t}=2$ for all $t\in\mathbb{N}$.
5.1.3 Results
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 4: The resulting SMCSs at four selected quantile levels with confidence level $\alpha=0.1$. Even though the SMCS behave differently at the different quantile levels, we may conclude that the ensemble models performed best as they are never excluded from the SMCSs. Second best is GT-deep, which only shows some deficiencies for $\tau=0.3$, and MOBS-gleam and the baseline, which seem to be approximately equally accurate. PSI-draft performs poorest and is excluded rather quickly at each quantile level.
Figure 4 shows the resulting SMCSs for four selected quantile levels with a confidence level of $\alpha=0.1$. As a first observation, we see that the power of the SMCS varies with the choice of the quantile level $\tau$: Whereas the SMCS detects four deficient models for $\tau=0.3$, it only has excluded two models by the end of the observation period for $\tau=0.7$. For $\tau=0.1$, the SMCS needs considerably more observations to reject the three models, which are excluded much faster at the median. As argued before, the lower power at the tails of the distribution is probably due to the factor $\max\left\{\tau,1-\tau\right\}$ in the bound (18). In Appendix F, we show that the effect of lower power at the tails even increases if we do not incorporate the quantile level in our betting scheme and that our proposed betting scheme is capable to mitigate this difficulty. Overall, the results align with our expectation that the two ensemble models perform best as they are included in the SMCS for the entire time period at all quantile levels. The superior performance of the two ensemble models was expected since they combine the forecasts of all other models and thus have access to the largest information set. Interestingly, the SMCS does not detect any significant difference between ensemble and CDC-ensemble, where the latter was supposed to improve the former by combining just the currently best, rather than all, forecasting models. The naive baseline performs surprisingly well and even outperforms the epidemic model PSI-draft, which is excluded earlier than baseline for all given values of $\tau$. We may conclude that PSI-draft is not competitive with the other models and should therefore not be considered as an accurate forecasting model for the pandemic. Finally, the second epidemic model MOBS-gleam performs better than PSI-draft but not as good as the neural network based GT-deep according to the SMCSs.
To conclude, we again highlight the key feature that SMCSs allow us to collect statistical evidence sequentially over time. Hence, if we were back in the pandemic, by using SMCSs, we could monitor the models performances from the very beginning on and would not have to wait until we have collected enough observations to perform a statistical test for predictive ability. Quickly, we would reject PSI-draft as an accurate model and not consider it in our analysis and policy-making anymore. Also MOBS-gleam would probably have been rejected as a competitive model by approximately the end of 2021, and so we would have been left with the choice between less and less models in the course of the pandemic. Clearly, in this extreme situation, it has already been common practice to assess the forecasting models on a regular basis. However, this practice may lack the theoretical justification, and our methodology provides statistically safe methods to do so.
5.2 Postprocessing of wind gust predictions
Currently, weather services use numerical weather prediction (NWP) models for predicting the future state of the atmosphere. These NWP models try to quantify the forecast uncertainty using ensemble predictions, where each ensemble member represents a different future scenario. However, the resulting ensemble forecasts fail to adequately quantify the forecast uncertainty, as they are subject to systematic biases and dispersion errors. Hence, the forecasts ought to be statistically postprocessed to generate accurate and reliable predictions (Gneiting and Raftery, 2005; Vannitsem et al., 2018).
Schulz and Lerch (2022) present a comprehensive comparison of multiple approaches for statistical postprocessing of ensemble forecasts for wind gusts. Here, we will build on their case study and compare the different forecasting models for probabilistic wind gust prediction on an extended time period by using SMCSs under the weak hypothesis.
5.2.1 Data
In this application, we use the data set of Primo et al. (2024), which is an extension of the original data in Schulz and Lerch (2022). The forecast data is based on an operational NWP model from the German weather service (Deutscher Wetterdienst, DWD) in the period from 08/12/2010 to 30/06/2023 with corresponding observations at 174 geographically diverse weather stations in Germany. The forecasts are initialized at 00 UTC and have a forecast horizon of up to 21 hours. Here, we consider only those forecasts with a lead time of 18h referring to 18 UTC and the $K=166$ stations with at least 2,600 observations.
Next to the ensemble forecasts (EPS), we include the eight postprocessing methods applied by Schulz and Lerch (2022) in our comparison. Statistical postprocessing methods are typically distributional regression models that use ensemble predictions from the NWP model as input data. The methods in our comparison can be divided into three groups: Basic techniques that only make use of the underlying wind gust ensemble predictions (EMOS, MBM, IDR), established machine learning methods for postprocessing that incorporate additional covariates (EMOS-GB, QRF), and neural network-based approaches (DRN, BQN, HEN). While the first two groups fit a separate model for each station, the neural networks estimate one locally adaptive model for all stations. Further, the models differ in the resulting forecast distributions, which range from parametric (EMOS, EMOS-GB, DRN) to semi- and non-parametric types (all others). The postprocessing models are trained on the period from 2010 to 2015 using the same configuration as in Schulz and Lerch (2022) and evaluated on the remaining period from 2016 to 2023.
To improve the forecast quality, NWP models are continuously developed further and frequently updated. Within the period from 2010 to 2023, the underlying NWP model has undergone several updates, of which three change the systematic errors of the ensemble forecasts drastically (on 22/03/2017, 16/05/2018 and 10/02/2021). These model updates present a challenge for statistical postprocessing methods, as the corrections learned on previous NWP model versions may not lead to the same improvements when applied to the current model version. In essence, the training data becomes less representative of the test data after each update. The methods considered here have been trained on data until the end of 2015, which is before the first major model update in 2017. Still, they are applied to ensemble forecasts that have been generated by another NWP model version, which has undergone up to three major updates. Thus, we expect the behavior both of the ensemble predictions and the postprocessed forecasts to change systematically over time. For further details on the data and postprocessing methods, we refer to Schulz and Lerch (2022) and Primo et al. (2024).
5.2.2 Methods and implementation details
We assess the $m=9$ different forecasting models by the CRPS and construct SMCSs for the weak hypothesis as proposed in Section 3.4. By predictability of the forecasts, the CRPS differences are conditionally bounded. More precisely, for any predictive cdfs $F_{1}$ and $F_{2}$, we have
$\lvert\textrm{CRPS}(F_{1},y)-\textrm{CRPS}(F_{2},y)\rvert\leq\max\{\lvert\textrm{CRPS}(F_{1},z)-\textrm{CRPS}(F_{2},z)\rvert\mid z\in\mathcal{X}\},\quad y\in\mathbb{R},$ (19)
[…]
For each station, we thus compute the crossing points of all issued forecasts at all days in order to obtain predictable bounds of the CRPS differences by (19). In a second step, we use the transformation of Section 3.5 to obtain uniformly bounded score differences. That is, $\lvert d_{ij,t}^{k}\rvert\leq C/2$ for the resulting CRPS differences at station $k\leq K$, for $C=2$, $i,j\leq m$ and $t\in\mathbb{N}$. We let $\lambda=(2C)^{-1}=1/4$, compare Section 3.5, and choose $\alpha=0.1$ as our confidence level. In order to draw inference over all stations simultaneously, we consider the average scaled CRPS differences ${d}_{ij,t}=\sum_{k=1}^{K}{d}^{k}_{ij,t}/K$, for $i,j\leq 9$ and $t\in\mathbb{N}$.
5.2.3 Results
First, we take a look at the SMCS with respect to the loss differences averaged over all stations, from which we draw conclusions regarding the performance of postprocessing methods that align with the results in Schulz and Lerch (2022). The left panel of Figure 5 displays the resulting SMCS and shows that the two neural network-based approaches DRN and BQN, which performed best in the former study, are included in the SMCS for the entire time period. The other models fall out of the SMCS over time, where the order coincides with that found in the former study, e.g., the underlying ensemble is omitted first. In this case, where we average the performances over all stations, the SMCS behaves as for the stronger hypotheses, that is, once we eliminate a model, we do not include it anymore.
Second, we consider the SMCS station-wise. The right panel of Figure 5 gives a summary, as it shows the number of stations for which a forecasting model is included in the corresponding SMCS. We find that the EPS is excluded from most SMCSs already within the first year, that the basic techniques EMOS, MBM and IDR are also omitted relatively fast, and that the machine learning approaches EMOS-GB and QRF fall out of SMCSs successively over time, until early 2021 when the last NWP model update occurred. Interestingly, the number of stations that include the HEN network approach increases after the last NWP model update, a behavior only detectable using the weak hypothesis. Finally, the BQN and DRN approach are included at almost all stations over the entire period, only with DRN being excluded over the last year at some stations.
Refer to captionRefer to caption
Figure 5: SMCS averaged over the loss differences over all stations (left) and the number of stations where the method is included in the SMCS for each method dependent on the time (right). The vertical black lines indicate the three major NWP model updates.
Taking a closer look at the individual stations, we observe different kinds of local behavior that is not visible when averaging over the loss differences. According to the findings, we broadly divide the stations into three categories: Regular (56 stations), dynamic (55) and fluctuating stations (55). Figure 6 shows one exemplary station for each of the three categories as well as the mean size of the station-wise SMCSs over time. The SMCS of a regular station behaves (almost) equally as for the stronger hypotheses, which is also seen in the plot of the mean SMCS size over the time period. The dynamic stations show a different pattern. As for the regular stations, some of the inferior models are excluded after a certain time period. However, in contrast to the regular stations, some methods that have been eliminated before are now included again coming closer to the end of the time period. This behavior becomes apparent when looking at the mean size of the SMCSs, which increases after the last update of the underlying NWP model that systematically changes the predictive performance of the methods. These cases are interesting, since any test which would be performed just at the end of the evaluation period (or at some given time point in the middle) would probably not detect the deficiencies of these models. Consider for example the SMCS at station 10044, Leuchtturm Kiel, displayed at the topright of Figure 6, where EMOS, IDR and HEN have been excluded from the SMCS for long periods and have become competitive again later on. Even though these three models perform not as strongly as the benchmarks DRN and BQN, the SMCS indicates that they are still better than the remaining models for this station. At last, we have the fluctuating stations, where at least one model is repeatedly excluded and included from the SMCS. It shows a kind of ”borderline”-behaviour of the method and indicates that it is exactly at the border of being worse than the other models. As for the other two groups, this behavior is reflected in the mean size of the SMCS, which is more oscillating for the fluctuating stations than the other two categories. In all cases, the sequential nature of the SMCS yields a deeper insight into the performances of the methods than tests for predictive ability which are issued at single time points only.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 6: SMCSs for one exemplary station in each of the three categories (topleft: normal; topright: dynamic; bottomleft: fluctuating) and the mean size of the SMCS averaged over the stations of each category dependent on the time (bottomright). Figure 8 in Appendix G shows the spatial distribution of all station categories over Germany.
6 Discussion
We introduced sequential model confidence sets and provided techniques to construct them with respect to three important notions of forecast superiority. Sequential model confidence sets address a highly relevant problem and allow to continuously monitor the performances of some given statistical models. They incorporate the uncertainty of the undergoing selection by being flexible in size, rely on minimal distributional assumptions, and achieve safe anytime-valid guarantees.
Concerning future work, there are various possible directions to investigate and extend the methods developed so far: Firstly, our construction is limited to (conditionally) bounded score differences. Important loss functions such as the quadratic score or the logarithmic score do not satisfy this requirement, and it would be useful to study e-processes and SMCSs for these more general loss functions, compare to Remark 3. Secondly, the simulation study shows that our methods are conservative. Since we have to adjust our processes for multiple testing, we may lose a considerable amount of evidence. More work is needed to reduce the gap in power between the raw and the adjusted e-processes, possibly by adaptive weighting schemes of the e-processes. Alternatively, one could study sequential model confidence sets with time-uniform FDR control as discussed in Appendix D. However, our impression is that there is not much gain in power by replacing the FWER with the less restrictive FDR. Finally, it would be interesting to examine how our results extend to important information criteria, and to apply sequential model confidence sets for sequential model selection.
Data and replication material
Code in R (R Core Team, 2024) for the case studies and replication material for the simulations are available at https://github.com/GGavrilos/SMCS.
Acknowledgments
We thank Lutz Dümbgen for valuable discussions and inputs. Sebastian Arnold and Johanna Ziegel gratefully acknowledge financial support from the Swiss National Science Foundation. Computations for the simulations and parts of the wind gust case study have been performed on UBELIX (https://ubelix.unibe.ch/), the HPC cluster of the University of Bern. Benedikt Schulz gratefully acknowledges funding within the project C5 “Dynamical feature-based ensemble postprocessing of wind gusts within European winter storms” of the Transregional Collaborative Research Center SFB/TRR 165 “Waves to Weather” funded by the German Research Foundation (DFG). We thank Cristina Primo, Reinhold Hess and Sebastian Trepte for providing the forecast and observation data, and Robert Redl for assistance in data handling.
Appendix A Proper scoring rules and consistent scoring functions
In this section, we review some important concepts for forecast evaluation. We assume that we are interested in a real-valued unknown quantity and let $\mathcal{P}(\mathbb{R})$ be the family of all probability distributions on $\mathbb{R}$.
A statistical functional is a map $T:\mathcal{P}(\mathbb{R})\to 2^{\mathbb{R}}$. Point-valued forecasts for functionals should be compared using consistent scoring functions (Gneiting, 2011). A measurable map $S:\mathbb{R}\times\mathbb{R}\rightarrow\bar{\mathbb{R}}$ is a consistent scoring function for the functional $T$ relative to the class $\mathcal{P}\subseteq\mathcal{P}(\mathbb{R})$ if for all $x\in\mathbb{R}$, $F\in\mathcal{P}$, the integral $\int S(x,y)\,\mathrm{d}F(y)$ exists and
$\int S(t,y)\,\mathrm{d}F(y)\leq\int S(x,y)\,\mathrm{d}F(y),\quad\textrm{for all }t\in T(F).$
The Brier score (or quadratic score) $\textrm{BS}(x,y)=(x-y)^{2}$, $x,y\in\mathbb{R}$, and the quantile score
$\textrm{QS}_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)(x-y),\quad x,y\in\mathbb{R},$ (20)
are main examples of consistent scoring functions for the mean functional, and for the quantile at level $\tau\in(0,1)$, respectively. More generally, Ehm et al. (2016) show that a scoring function $S$ is consistent for the quantile functional at level $\tau\in(0,1)$ if and only if it is of the form
$S_{\tau}(x,y)=(\mathbbm{1}{\{y\leq x\}}-\tau)\big{(}g(x)-g(y)\big{)},\quad x,y\in\mathbb{R},$
for some increasing function $g:\mathbb{R}\to\mathbb{R}$, where we obtain the quantile score at (20) for $g$ equals the identity, and where the result extends for the mean and the expectile.
Probabilistic forecasts incorporate the intrinsic uncertainty of future events and comprise of a probability measure on the possible outcomes. Probabilistic forecasts should be compared and evaluated using proper scoring rules (Gneiting and Raftery, 2007). A proper scoring rule is a function ${\rm S}:\mathcal{P}\times\mathbb{R}\mapsto\bar{\mathbb{R}}$ for some class $\mathcal{P}\subseteq\mathcal{P}(\mathbb{R})$ such that ${\rm S}(F,\cdot)$ is measurable for any $F\in\mathcal{P}$, the integral $\int{\rm S}(G,y)F(\,\mathrm{d}y)$ exists, and
$\int{\rm S}(F,y)\,\mathrm{d}F(y)\leq\int{\rm S}(G,y)\,\mathrm{d}F(y),\quad\textrm{for all }F,G\in\mathcal{P}.$
That is, for a proper scoring rule ${\rm S}$, the expected score $\mathbb{E}_{Y\sim F}{\rm S}(G,Y)$ is minimized over all distributional forecasts $G\in\mathcal{P}$, if we correctly forecast the true distribution of the random variable $Y$. The most commonly used proper scoring rule for the evaluation of real-valued outcomes is the continuous ranked probability score (CRPS; Matheson and Winkler (1976)), which is defined for all $F\in\mathcal{P}(\mathbb{R})$ and $y\in\mathbb{R}$ as
$\textrm{CRPS}(F,y)=\int(F(x)-\mathbbm{1}\{y\leq x\})^{2}\,\mathrm{d}x.$ (21)
The CRPS is a continuous extension of the Brier score for binary events and is popular across application areas and methodological communities; see, e.g., Gneiting et al. (2005), Hothorn et al. (2014) and Henzi et al. (2021).
Appendix B Conditionally bounded CRPS and quantile score differences
The following results show that CRPS and quantile score differences are conditionally bounded by the predictability of the forecasts.
Lemma B.1.
For $F_{1},F_{2}\in\mathcal{P}(\mathbb{R})$, the CRPS difference
$\delta(y)=\textrm{CRPS}(F_{1},y)-\textrm{CRPS}(F_{2},y)$
[…]
Proof.
By the integral representation of the CRPS given at (21),
$\delta(y)$ $=\int\big{(}F_{1}(x)-\mathbbm{1}\{y\leq x\}\big{)}^{2}-\big{(}F_{2}(x)-\mathbbm{1}\{y\leq x\}\big{)}^{2}\,\mathrm{d}x$
$=\int F_{1}(x)^{2}-F_{2}(x)^{2}\,\mathrm{d}x+2\int_{y}^{\infty}F_{2}(x)-F_{1}(x)\,\mathrm{d}x,$
where the first summand is independent of $y\in\mathbb{R}$. The function $x\mapsto F_{2}(x)-F_{1}(x)$ is bounded in $[-1,1]$, with limit 00 as $x\to\pm\infty$ and changes sign at the crossing points of $F_{1}$ and $F_{2}$. ∎
Lemma B.2.
For $x_{1},x_{2}\in\mathbb{R}$ and $\tau\in(0,1)$, the quantile loss difference
$\delta(y)=\textrm{QS}_{\tau}(x_{1},y)-\textrm{QS}_{\tau}(x_{2},y)$
is bounded in $y\in\mathbb{R}$ with
$(\tau-1)\lvert x_{1}-x_{2}\rvert\leq\delta(y)\leq\tau\lvert x_{1}-x_{2}\rvert.$
Proof.
Assume w.l.o.g. that $x_{1}\leq x_{2}$. Then, we have
$\delta(y)$ $=\big{(}\mathbbm{1}\{y\leq x_{1}\}-\tau\big{)}(x_{1}-y)-\big{(}\mathbbm{1}\{y\leq x_{2}\}-\tau\big{)}(x_{2}-y)$
$=y\mathbbm{1}\{x_{1}<y\leq x_{2}\}-\tau(x_{1}-x_{2})+x_{1}\mathbbm{1}\{y\leq x_{1}\}-x_{2}\mathbbm{1}\{y\leq x_{2}\}$
and the bounds follow directly by differentiating cases. ∎
Appendix C Sequential family-wise error rate and family-wise validity
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for some hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$.
Following Vovk and Wang (2021), we call $f$ a symmetric e-merging function if it is invariant with respect to permutations of its arguments and if $f(E_{i},i\in I)$ is an e-value for any family of e-values $\{E_{i}\mid i\in I\}$. The most important e-merging function is the arithmetic mean as it essentially dominates any symmetric e-merging function (Vovk and Wang, 2021, Proposition 3.1.). Let $f$ be any symmetric e-merging function. Then, by the closure principle (Vovk and Wang, 2021, Appendix A.3.), for any $t\in\mathbb{N}$,
$E^{\star}_{i,t}=\min\{f(E_{j,t},j\in I)\mid I\subseteq\{1,\dots,m\},i\in I\},\quad i=1,\dots,m,$ (22)
are family-wise valid e-variables for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$. That is, there exists a dominating family of e-variables $\{E^{\mathbb{Q}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$, with $E^{\mathbb{Q}}\geq E^{\star}_{i,t}$ for any $i=1,\dots,m$ and $\mathbb{Q}\in\mathcal{H}_{i}$.
If we adjust the e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ by (22) with respect to the same e-merging function $f$ for all $t\in\mathbb{N}$, then it follows directly that $(E^{\star}_{1,t})_{t\in\mathbb{N}},\dots,(E^{\star}_{m,t})_{t\in\mathbb{N}}$ are family-wise valid e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$. Importantly, the corresponding sequential test given at (8) control the family-wise error rate (FWER) uniformly in time, that is, for any $I\subseteq\{1,\dots,m\}$ and $\mathbb{Q}\in\bigcap_{i\in I}\mathcal{H}_{i}$,
$\textrm{FWER}_{I}(\mathbb{Q})=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i,t}^{\star}\geq 1/\alpha\textrm{ for some }i\in I\right)\leq\alpha,\quad\textrm{forall }\alpha\in(0,1).$ (23)
To verify (23), consider a dominating family of e-processes $\{(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$. Then, for any $\alpha\in(0,1)$, $I\subseteq\{1,\dots,m\}$ and $\mathbb{Q}\in\bigcap_{i\in I}\mathcal{H}_{i}$,
$\textrm{FWER}_{I}(\mathbb{Q})\leq\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{t}^{\mathbb{Q}}\geq 1/\alpha\right)\leq\alpha.$
Appendix D SMCSs by FDR control
For $i=1,\dots,m$, let $(\Psi_{i,t})_{t\in\mathbb{N}}$ be a sequential testing procedure for some hypothesis $\mathcal{H}_{i}\subseteq\mathfrak{B}(\Omega)$. We let $R_{t}=\sum_{i=1}^{m}\Psi_{i,t}$ be the number of all rejected hypotheses (discoveries) at $t\in\mathbb{N}$. For $\mathbb{Q}\in\bigcup_{i=1}^{m}\mathcal{H}_{i}$, we let $F_{t}(\mathbb{Q})=\sum_{i\in I_{\mathbb{Q}}}\Psi_{i,t}$ be the number of all true null hypotheses that are rejected at time $t$. The false discovery rate (FDR) at $t\in\mathbb{N}$ of the given testing procedures is defined as
$\textrm{FDR}_{t}=\sup_{\mathbb{Q}\in\bigcup_{i=1}^{m}\mathcal{H}_{i}}\mathbb{E}_{\mathbb{Q}}\left[\frac{F_{t}(\mathbb{Q})}{\max\{1,R_{t}\}}\right].$
We say that the sequential tests $(\Psi_{1,t})_{t\in\mathbb{N}},\dots,(\Psi_{m,t})_{t\in\mathbb{N}}$ time-uniformly control the false discovery rate at level $\alpha\in(0,1)$ if, for all (possibly infinite) stopping times $\tau$,
$\textrm{FDR}_{\tau}\leq\alpha.$ (24)
Wang and Ramdas (2022) study FDR control with e-values in a nonsequential setting. Looking carefully at the proof of their Theorem 2, we see that their arguments directly extend to e-processes to obtain sequential tests that satisfy (24) as explained in the following.
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$. Following Wang and Ramdas (2022), we let $E_{[1],t}\geq\dots\geq E_{[m],t}$ be the order statistics of $E_{1,t},\dots,E_{m,t}$ in decreasing order. For a given $\alpha\in(0,1)$, we define
$i^{\star}_{t}=\max\left\{i\in\{1,\dots,m\}\;\Big{|}\;\frac{iE_{[i],t}}{m}\geq\frac{1}{\alpha}\right\},\quad t\in\mathbb{N}.$
Then, we have a time-uniform FDR control at level $\alpha$ for the sequential tests
$\Psi_{i,t}=(\mathbbm{1}\{E_{i,t}\geq E_{[i_{t}^{\star}],t}\}),\quad i=1,\dots,m,t\in\mathbb{N}.$ (25)
As our construction for the strong and uniformly weak hypothesis builds on e-processes, we can use the sequential tests at (25) with respect to the e-processes given at (9). That is, at each $t\in\mathbb{N}$, we reject the $i^{\star}_{t}$ hypotheses with the largest values of the corresponding e-processes, and obtain the alternative SMCS
ˇ$\check{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<E_{[i_{t}^{\star}]\cdot,t}\right\}.$
Importantly, this sequence does not necessarily satisfy the originally imposed coverage requirement at (4) or (5) anymore. In contrast, it bounds the expected fraction of wrongly rejected models by the total number of rejected models at any random time by $\alpha$, that is, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$ and any random stopping time $\tau$, we have
$\mathbb{E}_{\mathbb{Q}}\left[\frac{\lvert\check{\mathcal{M}}^{c}_{\tau}\cap\mathcal{M}^{\star}_{\tau}\rvert}{\max\{1,\lvert\check{\mathcal{M}}_{\tau}^{c}\rvert\}}\right]\leq\alpha,$
where ˇ$\check{\mathcal{M}}^{c}_{\tau}=\{1,\dots,m\}\setminus\check{\mathcal{M}}_{\tau}$ and $(\mathcal{M}^{\star})_{t\in\mathbb{N}}$ is given by (1) or (2).
Appendix E Proof of Lemma 3.5
Proof.
For $t\in\mathbb{N}$, we have $C_{t,1-\alpha}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid\lVert f_{t}(\boldsymbol{X})\rVert_{1}\leq m(m-1)/\alpha\}$, for
$f_{t}:\mathbb{R}_{0}^{m\times m}\to\mathbb{R}_{0}^{m\times m},\quad\boldsymbol{X}\mapsto\left(f_{ij,t}(X_{ij})\right)_{i\neq j}=\left(\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij}-\lambda_{ij}tx-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\}\right)_{i\neq j},$
and where $\lVert\boldsymbol{X}\rVert_{1}=\sum_{i,j=1}^{m}\lvert X_{ij}\rvert$ for $\boldsymbol{X}\in\mathbb{R}^{m\times m}$. For any $i\neq j$, $f_{ij,t}$ is a nonnegative and nonincreasing convex function. Thus, the second claim follows immediately since $\boldsymbol{A}\leq\boldsymbol{B}$ implies $f_{t}(\boldsymbol{A})\geq f_{t}(\boldsymbol{B})$. Let $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$, then, for any $\lambda\in[0,1]$,
$\lVert f_{t}\big{(}\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\big{)}\rVert_{1}\leq\lVert\lambda f_{t}(\boldsymbol{A})+(1-\lambda)f_{t}(\boldsymbol{B})\rVert_{1}\leq\lambda\lVert f_{t}(\boldsymbol{A})\rVert_{1}+(1-\lambda)\lVert f_{t}(\boldsymbol{B})\rVert_{1}\leq\frac{m(m-1)}{\alpha},$
where the first inequality holds by convexity and the nonnegativity of $f_{t}$, the second inequality by the axioms of a norm, and the third inequality by the fact that $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$. ∎
Appendix F Additional material for the Covid case study
F.1 Documentation of the models
The first forecasting model baseline is a naive model created by the Covid-19 forecast hub team as a comparison benchmark. Baseline simply outputs the outcome of the most recent week as a weekly median forecast and forms a predictive distribution around this median prediction based on the past weekly incidences. For more information, we refer to the supplementary material of Cramer et al. (2022b). Secondly, we consider the model MOBS-gleam, a global epidemic and mobility structured meta-population model (GLEaM), developed by the MOBS lab. It consists of a population-mobility layer, which uses transportation data to model the population mobility though a stochastic process, and of a disease model, which classifies each person in a population as susceptible, latent, infectious, or recovered, and estimates the rate in which a susceptible individual contracts the disease. For more details, we refer to the original MOBS lab publications by Balcan et al. (2009) and Balcan et al. (2010). The third model PSI-draft was developed by the Predictive Science Inc. It is a SEIRX model and estimates the reproductive number $R_{0}$ by classifying each individual of a population into susceptible (S), exposed (E), infected (I), removed (R), or quarantined (X). The model stratifies the population by age and disease severity, and incorporates the social restrictions imposed over the course of the pandemic, see the GitHub page of the Covid-19 forecast hub (https://github.com/reichlab/covid19-forecast-hub/tree/master/data-processed/PSI-DRAFT) or the Zoltar forecast archive (https://zoltardata.com/model/254) for more information. Further, we consider the neural network based model GT-deep, which includes some auto-regressive terms to ensure temporal consistency across consecutive predictions. GT-deep was the first purely data-driven model to be included in the Covid-19 forecast hub, in the sense that it makes no epidemiological assumptions about the spread of the pandemic. More detailed information can be found in the original publication by Rodriguez et al. (2021). Finally, we consider two models that made their predictions based on previously issued forecasts from other models: Firstly, ensemble, formed by the Covid-19 forecast hub team, which combines the latest predictions of all models that submitted a full set of forecasts at all $23$ quantile values. Whereas ensemble outputted the average forecast of all individual models at each quantile level for the first few weeks of the pandemic, it switched to the corresponding median forecast later on. As discussed in Cramer et al. (2022b), ensemble was the model used by the CDC in its public communications about the pandemic. Finally, CDC-ensemble, developed jointly by the Covid-19 forecast hub and CDC, is very similar to ensemble, but instead of the median of all models (at each quantile value), it considers only the 10 best models, measured by the average weighted interval score over the 12 most recent weeks, see Cramer et al. (2022a).
F.2 Comparison to the naive betting approach
Figure 7 shows the resulting SMCSs with respect to the naive approach, where $K_{ij,t}=2$ for all $i,j=1,\dots,m$ and $t\in\mathbb{N}$, compare to Section 5.1.2.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: The resulting SMCSs with respect to the naive choice $K_{ij,t}=2$ for all $i,j\leq m$ and $t\in\mathbb{N}$ with confidence level $\alpha=0.1$. Whereas, there are no significant differences to the adaptive betting scheme proposed in Section 5.1.2 for $\tau=0.3,0.5,0.9$, the adaptive approach is clearly superior for $\tau=0.1$.
Appendix G Station categories for the wind gust case study
Figure 8: Spatial distribution of the station categories over Germany.